# Fine-grained Abnormality Prompt Learning for Zero-shot Anomaly Detection

### v1

@@ Abstract
Current zero-shot anomaly detection (ZSAD) methods show remarkable success in prompting large pre-trained vision-language models to detect anomalies in a target dataset without using any dataset-specific training or demonstration. However, these methods are often focused on crafting/learning prompts that capture only coarse-grained semantics of abnormality, e.g., high-level semantics like ‘damaged’, ‘imperfect’, or ‘defective’ on carpet. They therefore have limited capability in recognizing diverse abnormality details with distinctive visual appearance, e.g., specific defect types like color stains, cuts, holes, and threads on carpet. To address this limitation, we propose FAPrompt, a novel framework designed to learn Fine-grained Abnormality Prompts for more accurate ZSAD. To this end, we introduce a novel compound abnormality prompting module in FAPrompt to learn a set of complementary, decomposed abnormality prompts, where each abnormality prompt is formed by a compound of shared normal tokens and a few learnable abnormal tokens. On the other hand, the fine-grained abnormality patterns can be very different from one dataset to another. To enhance their cross-dataset generalization, we further introduce a data-dependent abnormality prior module that learns to derive abnormality features from each query/test image as a sample-wise abnormality prior to ground the abnormality prompts in a given target dataset. Comprehensive experiments conducted across 19 real-world datasets, covering both industrial defects and medical anomalies, demonstrate that FAPrompt substantially outperforms state-of-the-art methods by at least 3%-5% AUC/AP in both image- and pixel-level ZSAD tasks. Code is available at https://github.com/mala-lab/FAPrompt.
@@ 1 Introduction
Anomaly Detection (AD) is a critical task in computer vision, aiming to identify instances that deviate significantly from the majority of data. It has a wide range of real-world applications, e.g., industrial inspection and medical imaging analysis (Pang et al., 2021; Cao et al., 2024). Traditional AD methods focus on learning specialized detectors with large training samples. Consequently, these methods often rely on application-specific, carefully curated datasets to train a detection model, making them inapplicable for application scenarios where such data access is not possible due to data privacy issue, or where the test data significantly differs from the training set due to substantial distribution shifts arising from new deployment environments or other natural variations in datasets. Zero-shot AD (ZSAD), which aims at learning generalist models for detecting anomalies in a target dataset without using any dataset-specific training or demonstration, has been recently emerging as a promising approach to address this limitation of traditional AD approaches.
@@
In recent years, large pre-trained vision-language models (VLMs) such as CLIP (Radford et al., 2021) have demonstrated impressive zero/few-shot recognition capabilities across a broad range of vision tasks, including the ZSAD task (Chen et al., 2023b; Jeong et al., 2023; Deng et al., 2023; Zhou et al., 2024). To leverage VLMs for AD, the methods craft/learn text prompts to extract the textual semantic of normal/abnormal from VLMs for matching visual anomalies. These methods, such as WinCLIP (Jeong et al., 2023) and AnoVL (Deng et al., 2023), attempt to capture a range of abnormality semantics for better ZSAD by including a wide variety of pre-defined state-aware tokens (e.g., using ‘damaged’, ‘imperfect’, or ‘defective’ to depict defects on different objects like carpet) or domain-aware tokens (e.g., ‘industrial’, ‘manufacturing’, or ‘surface’) into the text prompts. Others (Zhou et al., 2024; 2022a; 2022b) employ learnable text prompts to extract more general-purpose features for representing the normal/abnormal class, such as AnomalyCLIP (Zhou et al., 2024). However, these approaches are focused primarily on coarse-grained semantics of abnormality and have limited capability in recognizing diverse abnormality details with distinctive visual appearance, e.g., specific defect types like color stains, cuts, holes, and threads on carpet, as shown in Fig. 1 Left, leading to failures in detecting anomalies that are different from the coarse-grained abnormal patterns. A recent approach AnomalyGPT (Gu et al., 2023) deals with this issue by using detailed text description of abnormal objects through an additional Large Language Model (LLM), but it requires the reference samples from the target data, which is a different task from ZSAD. It also heavily relies on costly human annotations for the detailed textual descriptions.
Figure 1: Left: FAPrompt vs. two related methods. Right: Their image-level ZSAD results in AUROC.
To tackle these issues, we propose a novel framework, namely FAPrompt, designed to learn Fine-grained Abnormality Prompts for more accurate ZSAD. In contrast to previous prompting methods, FAPrompt focuses on learning the prompts that can model diverse fine-grained abnormality semantics without requiring detailed human annotations or text descriptions, as illustrated in the bottom of Fig. 1 Left. To this end, in FAPrompt we introduce a novel Compound Abnormality Prompting module, namely CAP, to learn a set of complementary, decomposed abnormality prompts on top of a normal prompt, where each abnormality prompt is formed by a compound of the same tokens in the normal prompt and a few learnable abnormal tokens. The insight of this design is rooted from our observation that each abnormal pattern can be considered as some unexpected patterns overlaying on top of common normal patterns, e.g., color stains on normal texture of carpet. Such a compound prompting strategy enables the learning of different abnormality semantics easily while maintaining abnormality prompts in good proximity to the normal prompt. This helps avoid learning trivial abnormality prompts that are too far away from the normal prompt, lacking discriminability for distinguishing normal and abnormal samples.
On the other hand, the fine-grained abnormality patterns can be very different from one dataset to another. Thus, to achieve better cross-dataset generalization, the learned fine-grained abnormality prompts should be adaptive to any target testing datasets. We therefore further introduce a Data-dependent Abnormality Prior module, namely DAP, to enhance the cross-dataset generalizability of the abnormal tokens in CAP. It learns to derive abnormality features from each query/test image as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to a given target dataset. This interaction between CAP and DAP enables the learning of abnormality prompts that have fine-grained semantics and are adaptive to different testing datasets, enabling better ZSAD across a wide range of image AD datasets, as shown in Fig. 1 Right.
@@
  • We propose a novel ZSAD framework FAPrompt. Unlike existing methods that capture coarse-grained semantics of abnormality only, FAPrompt offers an effective approach for learning adaptive fine-grained abnormality semantics without any reliance on detailed human annotations or text descriptions.
  • To achieve this, we first introduce a novel Compound Abnormality Prompting module (CAP) in FAPrompt. It learns a small set of complementary, decomposed abnormality prompts on top of the normal prompt via a compounding normal-abnormal token design and an orthogonal constraint among the abnormality prompts.
  • We further introduce a Data-dependent Abnormality Prior module (DAP). It learns to select the most relevant abnormal features from anomaly images while refraining from normal images for adapting the fine-grained abnormalities learned in CAP to a given target dataset.
  • Comprehensive experiments on 19 diverse real-world industrial and medical image AD datasets show that FAPrompt significantly outperforms state-of-the-art ZSAD models by at least 3%-5% AUC/AP in both image- and pixel-level detection tasks.
@@ 2.1 Conventional Anomaly Detection
There have been different types of AD approaches introduced over the years. In particular, one-class classification methods (Tax & Duin, 2004; Yi & Yoon, 2020; Bergman & Hoshen, 2020; Chen et al., 2022; Ruff et al., 2020) aim to compactly describe normal data using support vectors. Reconstruction-based methods (Akcay et al., 2019; Schlegl et al., 2019; Zavrtanik et al., 2021b; Yan et al., 2021; Zaheer et al., 2020; Zavrtanik et al., 2021a; Park et al., 2020; Hou et al., 2021; Xiang et al., 2023; Liu et al., 2023; Yao et al., 2023b; a) train models to reconstruct normal images, with anomalies identified through higher reconstruction errors. Distance-based methods (Pang et al., 2018; Defard et al., 2021; Cohen & Hoshen, 2020; Roth et al., 2022) detect anomalies by measuring the distance between the test image and normal images. Knowledge distillation methods (Deng & Li, 2022; Bergmann et al., 2020; Salehi et al., 2021; Wang et al., 2021; Cao et al., 2023; Tien et al., 2023; Zhang et al., 2023) focus on distilling normal patterns from pre-trained models and detecting anomalies by comparing discrepancies between the distilled and original features. However, these methods often rely on application-specific datasets to train the detection model, limiting their applicability in real-world scenarios where data access is restricted due to privacy concerns, proprietary restrictions, or resource constraints. Also, these approaches tend to struggle when there is a significant difference between the distribution of the training and test data.
@@ 2.2 Zero-Shot Anomaly Detection
ZSAD has been made possible due to the development of large pre-trained foundation models, such as vision-language models (VLMs). CLIP (Radford et al., 2021) has been widely used as a VLM to enable ZSAD on visual data (Jeong et al., 2023; Zhou et al., 2024; Deng et al., 2023; Chen et al., 2023a). CLIP-AC adapts CLIP for ZSAD by using text prompts designed for the ImageNet dataset as in  (Radford et al., 2021). By using manually defined textual prompts specifically designed for industrial AD dataset, WinCLIP (Jeong et al., 2023) achieves better ZSAD performance compared to CLIP-AC, but it often does not generalize well to non-defect AD datasets. APRIL-GAN (Chen et al., 2023a) adapts CLIP to ZSAD through tuning some additional linear layers with annotated auxiliary AD data. AnoVL (Deng et al., 2023) introduces domain-aware textual prompts and test time adaptation in CLIP to enhance the ZSAD performance. AnomalyCLIP (Zhou et al., 2024) employs learnable, object-agnostic textual prompts to extract more general-purpose text features for the normal and abnormal classes. All these methods are focused on crafting/learning prompts that capture only coarse-grained semantics of abnormality, failing to detect anomalies that exhibit different patterns from these coarse abnormal patterns. There are a number of other studies leveraging CLIP for AD, but they are designed for empowering few-shot (Gu et al., 2023; Zhu & Pang, 2024) or conventional AD task (Joo et al., 2023; Wu et al., 2024a; b; c).
Figure 2: Overview of FAPrompt. It consists of two novel modules, including the Compound Abnormality Prompting (CAP) module and the Data-dependent Abnormality Prior (DAP) module detailed in the top-right and bottom-right corners respectively. CAP is devised to learn fine-grained abnormality semantics without relying on detailed human annotations or text descriptions, while DAP is designed to adaptively select the most abnormal features from each query/test image as a sample-wise abnormality prior to enhance the cross-dataset generalizability of the abnormality prompts in CAP.
@@ 3.1 Preliminaries
VLM Backbone. To enable accurate ZSAD, large pre-trained VLMs are typically required. Following existing approaches (Chen et al., 2023b; Deng et al., 2023; Jeong et al., 2023; Zhou et al., 2024), the pre-trained CLIP (Radford et al., 2021) is used in our study, which comprises a visual encoder $f_{v}(\cdot)$ and a text encoder $f_{t}(\cdot)$, where the visual and text representations are well-aligned through pre-training on web-scaled text-image pairs.
@@ 3.2 Overview of FAPrompt
In this work, we propose a ZSAD framework FAPrompt to learn adaptive fine-grained abnormality semantics without any reliance on detailed human annotations or text descriptions. Fig 2 illustrates the overall framework of FAPrompt that consists of two novel modules, including Compound Abnormality Prompting module (CAP) and Data-dependent Abnormality Prior module (DAP). To be more specific, the proposed CAP module is devised to specify the design of fine-grained abnormality prompts. The key characteristic of CAP is to obtain the abnormality prompts via a compound prompting method, where we have one normal prompt and multiple abnormality prompts are added on top of it. These normal and abnormal text prompts are then processed by the CLIP’s text encoder $f_{t}(\cdot)$ to generate the corresponding normal and abnormal text embeddings, respectively. For a given image $x$, FAPrompt extracts both an image token embedding $f_{v}(x)$ and a set of patch token embeddings $\mathbf{F}_{v}\in\mathbb{R}^{l\times d}$, with $l$ and $d$ respectively representing the length and dimensionality of $\mathbf{F}_{v}$. The prompts are then learned using $\mathcal{D}_{train}$ based on the similarity between the image and text embeddings, where the fine-grained abnormality prompts are aggregated into an abnormality prompt prototype before its use in similarity calculation. Further, the DAP module is introduced to improve cross-dataset generalization capability of the fine-grained abnormality prompts. DAP derives the most relevant abnormality features based on the given query/test image $x$, serving as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to the characteristics of a given target dataset. During training, the original parameters of CLIP remain frozen, and only the attached learnable tokens in the text encoder layers, along with the normal and fine-grained abnormality prompts, are optimized. Below we present these modules in detail.
@@ 3.3 Compound Abnormality Prompt Learning
Learning Fine-grained Abnormalities via Compound Normal and Abnormal Tokens. Previous approaches that rely on coarse-grained learnable text prompts fail to capture the fine-grained abnormality semantics for detecting diverse anomalies across different datasets. To address this, we propose the novel CAP module. The core insight is that abnormal samples typically exhibit different magnitude of deviation from their normal counterparts while still belonging to the same class. CAP models this by learning a set of complementary, decomposed abnormality prompts built on a shared normal prompt. Following previous work (Zhou et al., 2024), a set of learnable normal tokens and the fixed token ‘object’ are concatenated to define the normal text prompt $\mathcal{P}^{n}$. For the abnormality prompt, CAP aims to learn a small set of prompts of complementary semantics, denoted as $\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},...\mathcal{P}^{a_{K}}\}$, where each $\mathcal{P}^{a_{i}}$ is formed by a compound of the same tokens in the normal prompt $\mathcal{P}^{n}$ and a few learnable abnormal tokens. Formally, the normal and abnormality prompts can be defined as:
$\mathcal{P}^{n}=[V_{1}][V_{2}]...[V_{E}][object],$ (1)
$\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},...\mathcal{P}^{a_{K}}\},$
$with\ \mathcal{P}^{a_{i}}$ $=[V_{1}][V_{2}]...[V_{E}][A_{1}^{i}][A_{2}^{i}]...[A_{E^{\prime}}^{i}][object],$
@@
where $\{V_{1},V_{2},...V_{E}\}$ and $\{A_{1}^{i},A_{2}^{i},...A_{E^{\prime}}^{i}\}_{i=1}^{K}$ are learnable normal and abnormal tokens, respectively. This compound prompting strategy enables the learning of different abnormality semantics easily while maintaining abnormality prompts in good proximity to the normal prompt, supporting the learning of non-trivial, semantically-meaningful abnormality prompts.
@@
Learning Complementary Abnormality Prompts. To capture complementary fine-grained abnormalities and reduce redundant information captured by the abnormality prompts, it is essential to maximize the diversity among the fine-grained abnormalities. A straightforward approach would be to train distinct abnormal prompts on separate, annotated subsets with samples from different anomalous types. However, this would require extensive human annotations. To address this issue, we propose to add an orthogonal constraint loss $\mathcal{L}_{oc}$ into the abnormality prompts in CAP as a alternative method to encourage this diversity. Formally, the objective for this can be formulated as:
$\mathcal{L}_{oc}=\sum_{i,j\in K;i\neq j}abs(\frac{f_{t}(\mathcal{P}^{a_{i}})\cdot f_{t}(\mathcal{P}^{a_{j}})}{||f_{t}(\mathcal{P}^{a_{i}})||\times||f_{t}(\mathcal{P}^{a_{j}})||}),$ (2)
where the text encoder $f_{t}(\cdot)$ is used to extract the embeddings of the abnormality prompts, []delimited-[][$\cdot$][ ⋅ ] denotes inner product, $abs(\cdot)$ returns the absolute value, and |||||$\cdot$||| | ⋅ | | indicates the norm of vectors.
@@
To provide more representative embedding for the fine-grained abnormalities, we compute the prototype of the multiple abnormality prompt embeddings as the final fine-grained abnormality embedding via $\mathbf{F}_{a}=\frac{1}{|\mathcal{P}^{a}|}\sum_{\mathcal{P}^{a_{i}}\in\mathcal{P}^{a}}f_{t}(\mathcal{P}^{a_{i}})$. The normal text prompt embedding is $\mathbf{F}_{n}=f_{t}(\mathcal{P}^{n})$.
3.4 Learning to Select Data-dependent Abnormality Prior

### v2

@@ Abstract
Current zero-shot anomaly detection (ZSAD) methods show remarkable success in prompting large pre-trained vision-language models to detect anomalies in a target dataset without using any dataset-specific training or demonstration. However, these methods often focus on crafting/learning prompts that capture only coarse-grained semantics of abnormality, e.g., high-level semantics like ‘damaged’, ‘imperfect’, or ‘defective’ objects. They therefore have limited capability in recognizing diverse abnormality details that deviate from these general abnormal patterns in various ways. To address this limitation, we propose FAPrompt, a novel framework designed to learn Fine-grained Abnormality Prompts for accurate ZSAD. To this end, a novel Compound Abnormality Prompt learning (CAP) module is introduced in FAPrompt to learn a set of complementary, decomposed abnormality prompts, where abnormality prompts are enforced to model diverse abnormal patterns derived from the same normality semantic. On the other hand, the fine-grained abnormality patterns can be different from one dataset to another. To enhance the cross-dataset generalization, another novel module, namely Data-dependent Abnormality Prior learning (DAP), is introduced in FAPrompt to learn a sample-wise abnormality prior from abnormal features of each test image to dynamically adapt the abnormality prompts to individual test images. Comprehensive experiments on 19 real-world datasets, covering both industrial defects and medical anomalies, demonstrate that FAPrompt substantially outperforms state-of-the-art methods in both image- and pixel-level ZSAD tasks. Code is available at https://github.com/mala-lab/FAPrompt.
@@ 1 Introduction
Anomaly Detection (AD) is a critical task in computer vision, aiming to identify instances that deviate significantly from the majority of data. It has a wide range of real-world applications, e.g., industrial inspection and medical imaging analysis [37, 8]. Traditional AD methods often rely on application-specific, carefully curated datasets to train a specialized detection model, making them inapplicable for application scenarios where such data access is not possible due to data privacy issue, or where the test data significantly differs from the training set due to substantial distribution shifts arising from new deployment environments or other natural variations in datasets. To address this, Zero-shot Anomaly Detection (ZSAD) has emerged as a promising solution, enabling models to detect anomalies in unseen datasets without dataset-specific training.
Figure 1: Top: FAPrompt vs. previous methods. Bottom: 2-D t-SNE visualization of the prompt-wise anomaly score maps of three methods on samples from a BTAD dataset [35], with the maps generated based on the similarities between text prompt embeddings and image embeddings, as defined in Eq. 3.
@@
In recent years, large pre-trained vision-language models (VLMs) such as CLIP [39] have demonstrated impressive zero/few-shot recognition capabilities across a broad range of vision tasks. To leverage VLMs for ZSAD, existing methods [11, 25, 17, 66, 20] use craft/learn text prompts to extract the normal/abnormal textual semantics from VLMs for matching visual anomalies. Handcrafted text prompting methods such as WinCLIP [25] and AnoVL [17] attempt to capture a range of abnormality semantics for better ZSAD by using a wide variety of pre-defined textual description templates as text prompts. Others [66, 65, 64, 20] employ two learnable text prompts to extract more general-purpose features for distinguishing between normal and abnormal classes, such as AnomalyCLIP [66] (see Fig. 1 top-left).
However, existing methods often design/learn prompts taking inspiration from image classification, where the abnormality prompts are rudimentary when directly applied to the AD task, leading to a coarse-grained differentiation between normal and abnormal semantics, e.g., high-level semantic of a ‘damaged’, ‘imperfect’, or ‘defective’ carpet. Such prompt designs overlook the intrinsic relationships between normality and abnormality. In practice, abnormalities often manifest as subtle deviations from normal patterns rather than entirely distinct semantic categories, e.g., color stains on normal texture of carpet. As a result, the previous methods are prone to learn excessively simple and homogeneous abnormal patterns, as illustrated in Fig. 1 bottom (1), making it difficult to model diverse distributions of abnormal samples, thereby constraining their generalization across different datasets.
To tackle these issues, we propose a novel framework, namely FAPrompt, designed to learn Fine-grained Abnormality Prompts for accurate ZSAD. In contrast to previous prompting methods, FAPrompt focuses on learning the prompts that can model diverse, fine-grained abnormality semantics. To this end, in FAPrompt, we introduce a Compound Abnormality Prompt learning (CAP) module, which learns a set of complementary, decomposed abnormality prompts on top of a normal prompt, as illustrated in Fig. 1 top-right. This compound prompting method ensures that all abnormality prompts are semantically adherent to the normality captured in the normal prompt while introducing learnable deviations to capture fine-grained abnormalities. To enhance diversity and minimize redundancy among the abnormality prompts, an orthogonal constraint is further introduced to enforce that each abnormality prompt captures a unique aspect of abnormality while maintaining strong discriminability from the normal class.
On the other hand, most existing methods fail to leverage important prior information hidden in each query/test image, limiting the generalizability of text prompts and their adaptability to unseen datasets. CoCoOp [64] attempts to address this issue by incorporating instance-conditional information from image token embeddings. However, its design is ineffective for AD due to two reasons. i) It focuses on class-level semantic difference which can incorporate only coarse-grained semantics into the prompt learning. ii) It can introduce noise into abnormality prompts when the input images are normal samples since they lack abnormality-related information, reducing the discriminability of the learned abnormality prompts (see Table 4).
To overcome this challenge, we propose a Data-dependent Abnormality Prior learning (DAP) module. It learns to derive abnormality features from the most anomalous patches of each query/test image as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to individual input images. To avoid using noise (normal features) as the abnormality prior, we impose an abnormality prior learning loss to exclude features from the normal training samples.
The joint force of CAP and DAP enables FAPrompt to learn adherent yet diverse fine-grained semantics, as shown in Fig. 1 bottom (3). This effect cannot be achieved by simple ensemble methods that learn a set of abnormality prompts based on different base models, as they can learn redundant prompts, as shown in Fig. 1 bottom (2). Additional results and more t-SNE visualization with ensemble methods can be found in Sec. 4.2 and Appendix C.3.
@@
  • We propose a novel ZSAD framework FAPrompt. Unlike existing methods that capture coarse-grained semantics of abnormality only, FAPrompt offers an effective approach for learning diverse fine-grained abnormality semantics.
  • To achieve this, we introduce two novel modules – CAP and DAP in FAPrompt. CAP learns a small set of complementary, decomposed fine-grained abnormality prompts via a compounding normal-abnormal token design and an orthogonal constraint, whereas DAP aims to enhance cross-dataset generalization by learning sample-wise abnormality prior from the most abnormal features of abnormal training samples while refraining the prior being affected by normal training images.
  • Comprehensive experiments on 19 diverse real-world industrial and medical image AD datasets show that FAPrompt significantly outperforms state-of-the-art (SotA) ZSAD models by at least 3%-5% improvement in both image- and pixel-level detection tasks.
@@ 2.1 Conventional Anomaly Detection
There have been different types of AD approaches introduced over the years [37, 51, 8, 68]. In particular, one-class classification methods [47, 59, 3, 12, 41] aim to compactly describe normal data using support vectors. Reconstruction-based methods [1, 43, 62, 56, 60, 61, 38, 24, 55, 34, 58, 57] train models to reconstruct normal images, with anomalies identified through higher reconstruction errors. Distance-based methods [36, 14, 13, 40] detect anomalies by measuring the distance between the test image and normal patterns. Knowledge distillation methods [16, 5, 42, 49, 7, 48, 63] focus on distilling normal patterns from pre-trained models and detecting anomalies by comparing discrepancies between the distilled and original features. However, these methods often rely on application-specific datasets to train the detection model, limiting their applicability in real-world scenarios where data access is restricted due to privacy/resource issues, or where the distribution of test data significant differs from that of the training data.
@@ 2.2 Zero-Shot Anomaly Detection
ZSAD has been made possible due to the development of large pre-trained foundation models, such as vision-language models (VLMs). CLIP [39] has been widely used as a VLM to enable ZSAD on visual data [25, 66, 17, 10, 9]. CLIP-AC adapts CLIP for ZSAD by using text prompts designed for the ImageNet dataset as in  [39]. By using manually defined textual prompts specifically designed for industrial AD dataset, WinCLIP [25] achieves better ZSAD performance compared to CLIP-AC, but it often does not generalize well to non-defect AD datasets.
APRIL-GAN [10] adapts CLIP to ZSAD through tuning some additional linear layers with annotated auxiliary AD data. AnoVL [17] introduces domain-aware textual prompts and test time adaptation in CLIP to enhance the ZSAD performance. AnomalyCLIP [66], AdaCLIP [9] and FiLO [20] employ learnable textual/visual prompts to extract more general-purpose features for the normal and abnormal classes. All these methods are focused on crafting/learning prompts that capture only coarse-grained semantics of abnormality, failing to detect anomalies that exhibit different patterns from these coarse abnormal patterns. There are a number of other studies leveraging CLIP for AD, but they are designed for empowering few-shot [19, 67, 32] or conventional AD task [28, 52, 54, 53].
Figure 2: Overview of FAPrompt. It consists of two novel modules: the Compound Abnormality Prompt learning (CAP) (bottom-right) and the Data-dependent Abnormality Prior learning (DAP) (top-right). CAP is devised to learn fine-grained abnormality semantics without relying on detailed human annotations or text descriptions, while DAP is used to adaptively select the most abnormal features from each query/test image as a sample-wise abnormality prior to enhance the cross-dataset generalizability of the abnormality prompts in CAP.
@@ 3.1 Preliminaries
VLM Backbone. To enable accurate ZSAD, large pre-trained VLMs are typically required. Following existing approaches [17, 25, 66, 9], we select CLIP [39] in our study, which comprises a visual encoder $f_{v}(\cdot)$ and a text encoder $f_{t}(\cdot)$, where the visual and text features are well-aligned through pre-training on web-scaled text-image pairs.
@@ 3.2 Overview of FAPrompt
In this work, we propose a ZSAD framework FAPrompt to learn fine-grained abnormality semantics without any reliance on detailed human annotations or text descriptions. Fig 2 illustrates the overall framework of FAPrompt that consists of two novel modules, including Compound Abnormality Prompt learning (CAP) module and Data-dependent Abnormality Prior learning (DAP) module.
To be more specific, the proposed CAP module is devised to specify the design of fine-grained abnormality prompts. The key characteristic of CAP is to obtain the abnormality prompts via a compound prompting method, where we have one normal prompt and multiple abnormality prompts are added on top of it. These normal and abnormal text prompts are then processed by the CLIP’s text encoder $f_{t}(\cdot)$ to generate the corresponding normal and abnormal text embeddings, respectively. For a given image $x$, FAPrompt extracts both an image token embedding $f_{v}(x)$ and a set of patch token embeddings $\mathbf{F}_{v}\in\mathbb{R}^{l\times d}$, with $l$ and $d$ respectively representing the length and dimensionality of $\mathbf{F}_{v}$. The prompts are then learned using $\mathcal{D}_{train}$ based on the similarity between the visual and text embeddings, where the fine-grained abnormality prompts are aggregated into an abnormality prompt prototype before its use in similarity calculation.
Further, the DAP module is introduced to improve cross-dataset generalization capability of the fine-grained abnormality prompts. DAP derives the most relevant abnormality features from a given query/test image $x$, serving as a sample-wise abnormality prior to dynamically adapt the abnormality prompts in CAP to individual samples of a given target dataset. During training, the original parameters of CLIP remain frozen, and only the attached learnable tokens in the text encoder layers, along with the normal and fine-grained abnormality prompts, are optimized. Below we present these modules in detail.
@@ 3.3 Compound Abnormality Prompt Learning
Learning Fine-grained Abnormalities via Compound Normal and Abnormal Tokens. Previous approaches that rely on coarse-grained learnable text prompts fail to capture the fine-grained abnormality semantics for detecting diverse anomalies across different datasets. To address this, we propose the novel CAP module. The core insight is that abnormal samples typically exhibit different magnitude of deviation from their normal counterparts while still being adherent to the normal class. CAP models this by learning a set of complementary, decomposed abnormality prompts built on a shared normal prompt. To be specific, a set of learnable normal tokens and the fixed token ‘object’ are concatenated to define the normal text prompt $\mathcal{P}^{n}$. For the abnormality prompt, CAP aims to learn a small set of prompts of complementary semantics, denoted as $\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},...\mathcal{P}^{a_{K}}\}$, where each $\mathcal{P}^{a_{i}}$ is formed by a compound of the same tokens in the normal prompt $\mathcal{P}^{n}$ and a few learnable abnormal tokens. Formally, the normal and abnormality prompts can be defined as:
$\mathcal{P}^{n}=[V_{1}][V_{2}].[V_{E}][object],$ (1)
$\mathcal{P}^{a}=\{\mathcal{P}^{a_{1}},\mathcal{P}^{a_{2}},.\mathcal{P}^{a_{K}}\},$
$with\ \mathcal{P}^{a_{i}}$ $=[V_{1}][V_{2}].[V_{E}][A_{1}^{i}][A_{2}^{i}].[A_{E^{\prime}}^{i}][object],$
@@
where $\{V_{1},V_{2},...V_{E}\}$ and $\{A_{1}^{i},A_{2}^{i},...A_{E^{\prime}}^{i}\}_{i=1}^{K}$ are learnable normal and abnormal tokens, respectively. This compound prompting strategy enables the learning of different abnormality semantics easily while maintaining abnormality prompts being adherent to the normality in the normal prompt, supporting the learning of semantically-meaningful abnormality prompts.
@@
Learning Complementary Abnormality Prompts. To capture complementary fine-grained abnormalities and reduce redundant information captured by the abnormality prompts, it is essential to maximize the diversity among the fine-grained abnormalities. A straightforward approach would be to train distinct abnormality prompts on separate, annotated subsets with samples from different anomalous types. However, this would require extensive human annotations. To address this issue, we propose to add an orthogonal constraint loss $\mathcal{L}_{oc}$ into the abnormality prompts in CAP as a alternative method to encourage this diversity. Formally, the objective for this can be formulated as:
[…]
@@
where the text encoder $f_{t}(\cdot)$ is used to extract the embeddings of the abnormality prompts, $\cdot$] denotes inner product, $abs(\cdot)$ returns the absolute value, and $||\cdot||$ indicates the norm of vectors. To provide more representative embedding for the fine-grained abnormalities, we compute the prototype of the multiple abnormality prompt embeddings as the final fine-grained abnormality embedding via $\mathbf{F}_{a}=\frac{1}{|\mathcal{P}^{a}|}\sum_{\mathcal{P}^{a_{i}}\in\mathcal{P}^{a}}f_{t}(\mathcal{P}^{a_{i}})$. The normal text prompt embedding is $\mathbf{F}_{n}=f_{t}(\mathcal{P}^{n})$.
3.4 Data-dependent Abnormality Prior Learning